# Exponential separations between classical and quantum learners

### v1

@@ 3.3 A learning separation with efficiently evaluatable concepts
3.4 A learning separation with a fixed hypothesis class
@@
Fix an $n$-bit integer $N=pq$, where $p$ and $q$ are two $\lfloor n/2\rfloor$-bit primes such that $\gcd\big{(}3,(p-1)(q-1)\big{)}=1$. We define the cube root identification concept class as $\mathcal{C}_{n}^{\mathrm{DCRI}}=\{c_{m}\}_{m\in\mathbb{Z}^{*}_{N}}$, with
$c_{m}(x)=\mathrm{bin}(m^{3}\text{ mod }N,\text{ }\mathrm{int}(x_{1}:\dots:x_{\lfloor\log n\rfloor})),$
where $\mathrm{bin}(y,k)$ denotes the $k$th bit of the binary representation of $y$, and $\mathrm{int}(x_{1}:\dots:x_{\lfloor\log n\rfloor})$ is the integer encoded by the first $\lfloor\log n\rfloor$-bits of $x\in\{0,1\}^{n}$.
We show that the cube root identification concept class is not classically learnable with a fixed hypothesis class under the Discrete Cube Root Assumption (DCRA) discussed in Section 3.2. To show that the existence of an efficient classical learner violates the DCRA, we let $e\in\mathbb{Z}^{*}_{N}$ and show we how an efficient classical learner can efficiently compute $m=f_{N}^{-1}(e)$. First, we generate examples $(x,\mathrm{bin}(e,k))$, where $k=\mathrm{int}(x_{1}:\dots:x_{\lfloor\log n\rfloor})$. When plugging these examples into an efficient classical learner it will with high probability identify an $m^{\prime}$ such that $(m^{\prime})^{3}\equiv m\mod N$. Since $x\mapsto x^{3}\mod N$ is a bijection on $\mathbb{Z}^{*}_{N}$ we find that $m=m^{\prime}$, and thus conclude that an efficient classical learner can indeed efficiently compute the solution to our discrete cube root instance.
To establish that the cube root identification concept class is quantumly proper learnable, we first note that using $\mathcal{O}(\mathrm{poly}(n))$ examples of a concept $c_{m}$ under the uniform distribution we can with high probability reconstruct the full binary representation of $m^{3}$. Next, since $N$ is known we can use Shor’s algorithm [Sho99] to compute $d$ such that $(m^{3})^{d}\equiv m\text{ mod }N$, which allows us to correctly identify the concept $c_{m}$. Note that the quantum learner needs access to the data to obtain a full reconstruction of the binary representation of $m^{3}$. We summarize the learning separation of the cube root identification concept class in Theorem 4, and we defer the proof to Appendix C.
@@
$L_{\mathrm{DCRI}}=\big{(}\{\mathcal{C}_{n}^{\mathrm{DCRI}}\}_{n\in\mathbb{N}},\{\mathcal{D}^{U}_{n}\}_{n\in\mathbb{N}})$ exhibits a $\mathsf{C}_{\mathcal{H}}/\mathsf{Q}_{\mathcal{H}}$ separation, where $\mathcal{H}=\mathcal{C}^{\mathrm{DCRI}}$ and $\mathcal{D}^{U}_{n}$ denotes the uniform distribution over $\mathbb{Z}_{N}^{*}$.
In conclusion, the cube root identification concept class exhibits a separation in the setting where the learner is constrained to only output hypotheses from a fixed hypothesis class. In fact, this is a separation in the proper efficient PAC learning framework, since the hypothesis class is the same as the concept class. Since in this setting it is not required to evaluate the concepts on unseen examples, the classical hardness has to lie in identifying rather than evaluating the concept generating the examples. We remark that for the cube root identification concept class it is not possible to obtain a similar learning separation for a singleton concept class where a learner does not require any data (see also Observation 1).
@@ 4 Learning separations without efficient data generation
In the quantum machine learning community there is an often-mentioned conjecture that quantum machine learning is most likely to have its advantages for data that is generated by a “genuine quantum process” [^1] Recently, there have been notable developments that have yielded contrasting conclusions. For instance, in [${}^{+}$22b], surprisingly complex physics problems are efficiently learned by classical learners. We will briefly discuss this in Section 5.1.. We understand this to mean that the concepts generating the data are $\mathsf{BQP}$-complete or perhaps $\mathsf{DQC1}$-complete. It is worth noting that if concepts in $\mathsf{BQP}$ or $\mathsf{DQC1}$ that are not in $\mathsf{BPP}$ are already considered a “genuine quantum process”, then the discrete logarithm concept class discussed in Section 3.1 suffices. However, we aim to investigate learning separations beyond these concepts, i.e., where the concepts are $\mathsf{BQP}$-complete.
@@
As discussed in Section 2.2, assuming that the concepts are not in $\mathsf{HeurP/poly}$ is sufficient to ensure that the concept class is not classically learnable. Intuitively, this is because if the concepts were classically learnable, the examples could be used to construct an advice string that, together with an efficient classical learning algorithm, would put the concepts in $\mathsf{HeurP/poly}$. By combining this with our approach to ensure quantum learnability, we can show that if there exists a family of polynomially-sized concept classes consisting of $\mathsf{BQP}$-complete concepts that are not in $\mathsf{HeurP/poly}$, then this family of concept classes exhibits a $\mathsf{CC/QQ}$ separation. Moreover, in Section 4.1 we discuss how several of these separations can be build around data that is generated by a “genuine quantum process”. The following theorem summarizes our findings, and we defer the proof to Appendix D.
@@
At face value, it may not be clear whether there exist concept classes that satisfy both conditions (a) and (c), since condition (a) puts the concepts in $\mathsf{BQP}$ and it may not be clear how large $\mathsf{HeurP/poly}$ is relative to $\mathsf{BQP}$. Notably, it is known that if the discrete logarithm is not in $\mathsf{BPP}$, then it is also not in $\mathsf{HeurP}$ under certain distributions. Additionally, it is widely believed that a polynomial amount of advice does not significantly improve the computational complexity of solving the discrete logarithm problem. Hence, it is plausible to imagine the existence of problems $L\in\mathsf{BQP}$ for which there is a distribution $\mathcal{D}$ such that $(L,\mathcal{D})\not\in\mathsf{HeurP/poly}$. Moreover, it is interesting to observe that if there exists a single $L\in\mathsf{BQP}$ that is not in $\mathsf{HeurP/poly}$ under some distribution, then for every $\mathsf{BQP}$-complete problem there exists a distribution under which it is not in $\mathsf{HeurP/poly}$. We summarize this in the lemma below, and we defer the proof to Appendix D.1.
@@
In summary, to obtain a learning separation for data generated by a “genuine quantum process”, it is sufficient to have a single problem $L\in\mathsf{BQP}$ that lies outside $\mathsf{HeurP/poly}$ under some distribution. An example of such a problem is the discrete logarithm. However, the resulting distribution under which the $\mathsf{BQP}$-complete problem lies outside of $\mathsf{HeurP/poly}$ is artificial as it comes from explicitly encoding the discrete logarithm into the learning problem through the reduction to the $\mathsf{BQP}$-complete problem. Besides the discrete logarithm, little is known about the heuristic hardness of problems in $\mathsf{BQP}$ (especially those that are considered “genuinely quantum”). Therefore, the question arises as to what additional properties are required for a $\mathsf{BQP}$-complete problem to lie outside $\mathsf{HeurP/poly}$ under some distribution. We show that a worst-case to average-case reduction combined with the assumption that $\mathsf{BQP}\not\subseteq\mathsf{P/poly}$ is sufficient for this purpose. While the question of $\mathsf{BQP}\not\subseteq\mathsf{P/poly}$ remains open, we proceed under this assumption based on its implications for cryptography. Specifically, if $\mathsf{BQP}\subseteq\mathsf{P/poly}$, then problems like the discrete logarithm would be in $\mathsf{P/poly}$, which would break cryptographic systems assumed to be secure [^3] In cryptography it is common to assume non-uniform adversaries (i.e., with computational resources of $\mathsf{P/poly}$), and even in this case most public-key cryptosystems such as RSA and Diffie-Hellman are still assumed to be safe).. Under the assumption that $\mathsf{BQP}\not\subseteq\mathsf{P/poly}$, the only missing piece is that our problem $L\in\mathsf{BQP}$ that lies outside $\mathsf{P/poly}$ is random self-reducible with respect to some distribution (i.e., it admits a worst-case to average case reduction as discussed in Section 2.2). We summarize these findings in the lemma below, and we defer the proof to Appendix D.2.
@@ 5 Connections to other works on (quantum) learning tasks
In this section we discuss other topics of relevance. First, in Section 5.1, we discuss the implications and limitations of the milestone work of Huang et al. [${}^{+}$22b] on establishing learning separations from physical systems. Next, in Section 5.2, we discuss how having access to data radically enhances what can be efficiently evaluated by discussing the example of evaluating parameterized quantum circuits. Afterwards, in Section 5.3, we discuss how two physically-motivated problems (i.e., Hamiltonian learning, and identifying order parameters for phases of matter) fit in the PAC learning setting where the learner is constrained to output hypotheses from a fixed hypothesis class.
@@ 5.1 Provably efficient machine learning with classical shadows
In the milestone work of Huang et al. [${}^{+}$22b], the authors design classical machine learning methods (in part built around the classical shadow paradigm) that can efficiently learn quantum many-body problems. One of the problems studied in [${}^{+}$22b] is that of predicting ground states of Hamiltonian. More precisely, for a family of Hamiltonians $H(x)$ with ground states $\rho_{H}(x)$, one wants to predict the expectation value of some observable $O$ when measured on $\rho_{H}(x)$. That is, one wants to efficiently learn to evaluate the function
@@
One of the main things that [${}^{+}$22b] show is that given a polynomial number of data points, one is able to efficiently evaluate the functions in Eq. (15) with a constant expected error under certain criteria. Recall that in Section 4.1 we argued that concepts based on physical systems can be used as a source of learning separations. Since these concepts are of a similar form as the functions described in Eq. (15), one might wonder how the results of Huang et al. relate.
Let us take a closer look at the requirements of the methods described in [${}^{+}$22b]. Firstly, the Hamiltonians $H(x)$ must all be geometrically-local, and the observable $O$ must be a sum of polynomially many local observables $O=\sum_{i=1}^{L}O_{i}$ such that $\sum_{i=1}^{L}||O_{i}||$ is bounded by a constant. Additionally, the Hamiltonians $H(x)$ must all have a constant spectral gap (i.e., the difference between the smallest and the next smallest eigenvalue) and they must depend smoothly on $x$ (or more precisely, the average gradient of the function in Eq. (15) must be bounded by a constant). One might wonder what will happen if we relax the above requirements, while simultaneously maintaining the fact that a quantum computer would still be able to evaluate the function in Eq. (15) (and hence build a learning separation around it based on Theorem 5).
Two possible relaxations of the requirements are the absence of a constant spectral gap (while maintaining an inverse polynomial spectral gap) and a reduced smoothness dependency of the Hamiltonian family on $x$ (i.e., compared to what is required for the methods of [${}^{+}$22b]). It turns out that if one relaxes these requirements, then under cryptographic assumptions the methods proposed by Huang et al. are no longer capable of evaluating the function in Eq. (15) with constant expected error. More precisely, any classical machine learning method that would still be able to evaluate the function in Eq. (15) up to constant expected error under the relaxed assumptions would be able to solve $\mathsf{DLP}$ in $\mathsf{P/poly}$, which contradicts certain cryptographic assumptions. We provide a formal statement of this in the following theorem, the proof of which is deferred to Appendix E.
@@
In conclusion, Theorem 6 shows that any method similar to that of [${}^{+}$22b] cannot learn to predict ground state properties of certain physical systems discussed in Section 4.1. Moreover, there are a few subtle differences between the setup of [${}^{+}$22b] and the one discussed in this paper. Firstly, the classical shadow paradigm uses data that is different from the PAC learning setting (i.e., the data does not correspond to evaluations of the function it aims to predict). This distinction in setup makes the approach of [${}^{+}$22b] more versatile, as their data can be utilized to evaluate multiple different observables (moreover, their methods also work in the PAC setting). Secondly, the functions $f_{H,O}$ in Eq. (15) are real-valued, which differs from our paper where we investigate functions that map onto a discrete label space. It is possible to address this difference by applying a threshold function to $f_{H,O}$ after it is learned. However, this thresholding introduces a mismatch in the types of data, as it would involve using real-valued data to learn a function with discrete values (which is clearly different from the PAC setting).

### v2

@@ Conclusion
3.4 Learning separation with a fixed hypothesis class
@@
We define the cube root identification concept class as
$\mathcal{C}^{\mathrm{DCRI}}_{n}=\Big{\{}c_{(N,m)}\mid N=pq\text{, $p$ and $q$ two $\lfloor n/2\rfloor$-bit primes s.t.\ $\gcd\big{(}3,(p-1)(q-1)\big{)}=1$},\text{ }m\in\mathbb{Z}_{N}^{*}\Big{\}},$
where
$c_{(N,m)}(b,x)=\begin{cases}\mathrm{bin}(m^{3}\text{ mod }N,\text{ }\mathrm{int}(x_{1}:\dots:x_{\lfloor\log n\rfloor}))\quad&\text{if }b=0\\ \mathrm{bin}(N,\mathrm{int}(x_{1}:\dots:x_{\lfloor\log n\rfloor}))\quad&\text{if }b=1.\end{cases}$
for $b\in\{0,1\}$, $x\in\{0,1\}^{n}$.
Remark.
Here $\mathrm{bin}(y,k)$ denotes the $k$th bit of the binary representation of $y$, and we additionally use $\mathrm{int}(x_{1}:\dots:x_{\lfloor\log n\rfloor})$ to denote the integer encoded by the first $\lfloor\log n\rfloor$-bits of $x\in\{0,1\}^{n}$.
We summarize the learning separation of the cube root identification concepts in Theorem 7, we defer the proof to Appendix E, and we discuss various aspects of this learning separation below.
@@
$L_{\mathrm{DCRI}}=\left(\{\mathcal{C}^{\mathrm{DCRI}}_{n}\}_{n\in\mathbb{N}},\{\mathcal{D}^{U}_{n}\}_{n\in\mathbb{N}}\right)$ exhibits a $\mathsf{C}_{\mathcal{H}}/\mathsf{Q}_{\mathcal{H}}$ separation assuming the intractability of the discrete cube root as outlined in [KV94b], where $\mathcal{D}_{n}^{U}$ denotes the uniform distribution over $\mathcal{X}_{n}=\{0,1\}\times\{0,1\}^{n}$.
Quantum learnability
To establish that the cube root identification concept class is quantumly proper learnable, we first note that using $\mathcal{O}(\mathrm{poly}(n))$ examples of a concept $c_{m}$ under the uniform distribution we can with high probability reconstruct the full binary representation of $m^{3}$. Since we can also with high probability recover $N$ using the same amount of examples, we can use Shor’s algorithm [Sho99] to compute $d$ such that $(m^{3})^{d}\equiv m\text{ mod }N$ (allowing us to correctly identify the concept $c_{m}$). We remark that the quantum learner needs access to the data in order to obtain a full reconstruction of the binary representation of both $m^{3}\mod N$ as well as the modulus $N$.
Hardness assumption
We show that the cube root identification concept class is not classically learnable with a fixed hypothesis class under the Discrete Cube Root Assumption (DCRA) discussed in Section 3.2. To show that the existence of an efficient classical learner violates the DCRA, we let $e\in\mathbb{Z}^{*}_{N}$ and show we how an efficient classical learner can efficiently compute $m=f_{N}^{-1}(e)$. First, we generate examples $(x,\mathrm{bin}(e,k))$ and $(x,\mathrm{bin}(N,k))$, where $k=\mathrm{int}(x_{1}:\dots:x_{\lfloor\log n\rfloor})$. When plugging these examples into an efficient classical learner it will with high probability identify an $m^{\prime}$ such that $(m^{\prime})^{3}\equiv m\mod N$. Since $x\mapsto x^{3}\mod N$ is a bijection on $\mathbb{Z}^{*}_{N}$ we find that $m=m^{\prime}$, and conclude that an efficient classical learner can efficiently compute the discrete cube root $f_{N}^{-1}(e)$.
Conclusion
The cube root identification concept class exhibits a separation in the setting where the learner is constrained to only output hypotheses from a fixed hypothesis class. In fact, this is a separation in the so-called proper PAC framework, since the hypothesis class is the same as the concept class. Since in this setting it is not required to evaluate the concepts on unseen examples, the classical hardness has to lie in identifying rather than evaluating the concept generating the examples. Note that for the cube root identification concept class it is not possible to obtain a learning separation for a singleton concept class where a learner does not require any data (see Observation 1).
@@ 4 Learning separations without efficient data generation
In the quantum machine learning community there is an often-mentioned conjecture that quantum machine learning is most likely to have its advantages for data that is generated by a “genuine quantum process” [^1] Recently, there have been notable developments that have yielded contrasting conclusions. For instance, in [HKT+22b], surprisingly complex physics problems are efficiently learned by classical learners. We will briefly discuss this in Section 5.1.. We understand this to mean that the concepts generating the data are $\mathsf{BQP}$-complete or perhaps $\mathsf{DQC1}$-complete. It is worth noting that if concepts in $\mathsf{BQP}$ or $\mathsf{DQC1}$ that are not in $\mathsf{BPP}$ are already considered a “genuine quantum process”, then the discrete logarithm concept class discussed in Section 3.1 suffices. However, we aim to investigate learning separations beyond these concepts, i.e., where the concepts are $\mathsf{BQP}$-complete.
@@
As discussed in Section 2.2, assuming that the concepts are not in $\mathsf{HeurP/poly}$ is sufficient to ensure that the concept class is not classically learnable. Intuitively, this is because if the concepts were classically learnable, the examples could be used to construct an advice string that, together with an efficient classical learning algorithm, would put the concepts in $\mathsf{HeurP/poly}$. By combining this with our approach to ensure quantum learnability, we can show that if there exists a family of polynomially-sized concept classes consisting of $\mathsf{BQP}$-complete concepts that are not in $\mathsf{HeurP/poly}$, then this family of concept classes exhibits a $\mathsf{CC/QQ}$ separation. Moreover, in Section 4.1 we discuss how several of these separations can be build around data that is generated by a “genuine quantum process”. The following theorem summarizes our findings, and we defer the proof to Appendix F.
@@
At face value, it may not be clear whether there exist concept classes that satisfy both conditions (a) and (c), since condition (a) puts the concepts in $\mathsf{BQP}$ and it may not be clear how large $\mathsf{HeurP/poly}$ is relative to $\mathsf{BQP}$. Notably, it is known that if the discrete logarithm is not in $\mathsf{BPP}$, then it is also not in $\mathsf{HeurBPP}$ under certain distributions. Additionally, it is widely believed that a polynomial amount of advice does not significantly improve the computational complexity of solving the discrete logarithm problem [CGK18]. Hence, it is plausible to imagine the existence of problems $L\in\mathsf{BQP}$ for which there is a distribution $\mathcal{D}$ such that $(L,\mathcal{D})\not\in\mathsf{HeurP/poly}$. Moreover, it is interesting to observe that if there exists a single $L\in\mathsf{BQP}$ that is not in $\mathsf{HeurP/poly}$ under some distribution, then for every $\mathsf{BQP}$-complete problem there exists a distribution under which it is not in $\mathsf{HeurP/poly}$. We summarize this in the lemma below, and we defer the proof to Appendix F.1.
@@
In summary, to obtain a learning separation for data generated by a “genuine quantum process”, it is sufficient to have a single problem $L\in\mathsf{BQP}$ that lies outside $\mathsf{HeurP/poly}$ under some distribution. An example of such a problem is the discrete logarithm. However, the resulting distribution under which the $\mathsf{BQP}$-complete problem lies outside of $\mathsf{HeurP/poly}$ is artificial as it comes from explicitly encoding the discrete logarithm into the learning problem through the reduction to the $\mathsf{BQP}$-complete problem. Besides the discrete logarithm, little is known about the heuristic hardness of problems in $\mathsf{BQP}$ (especially those that are considered “genuinely quantum”). Therefore, the question arises as to what additional properties are required for a $\mathsf{BQP}$-complete problem to lie outside $\mathsf{HeurP/poly}$ under some distribution. We show that a worst-case to average-case reduction combined with the assumption that $\mathsf{BQP}\not\subseteq\mathsf{P/poly}$ is sufficient for this purpose. While the question of $\mathsf{BQP}\not\subseteq\mathsf{P/poly}$ remains open, we proceed under this assumption based on its implications for cryptography. Specifically, if $\mathsf{BQP}\subseteq\mathsf{P/poly}$, then problems like the discrete logarithm would be in $\mathsf{P/poly}$, which would break cryptographic systems assumed to be secure [^3] In cryptography it is common to assume non-uniform adversaries (i.e., with computational resources of $\mathsf{P/poly}$), and even in this case most public-key cryptosystems such as RSA and Diffie-Hellman are still assumed to be safe).. Under the assumption that $\mathsf{BQP}\not\subseteq\mathsf{P/poly}$, the only missing piece is that our problem $L\in\mathsf{BQP}$ that lies outside $\mathsf{P/poly}$ is random self-reducible with respect to some distribution (i.e., it admits a worst-case to average case reduction as discussed in Section 2.2). We summarize these findings in the lemma below, and we defer the proof to Appendix F.2.
@@
An interesting observation can be made when we consider the constructions underlying Corollary 1 and let $L=\mathsf{DLP}$ to be the language corresponding to the discrete logarithm problem. Firstly, we note that $\mathsf{DLP}\in\mathsf{BQP}$ due to Shor’s algorithm [Sho99]. Secondly, since $\mathsf{DLP}$ is both random self-reducible [BM84] and random verifiable (i.e., we can generate examples under the uniform distribution as in Section 3.1), we know that $\mathsf{DLP}\in\mathsf{HeurBPP/samp}$ is equivalent to $\mathsf{DLP}\in\mathsf{BPP}$. By combining these observations with the constructions underlying Corollary 1 we obtain the lemma below, which shows that for every $\mathsf{BQP}$-$\mathsf{complete}$ problem we can construct a distribution $\mathsf{D}_{L}$ under which it exhibits a $\mathsf{CC}/\mathsf{QQ}$ separation unless $\mathsf{DLP}\in\mathsf{BPP}$ (we defer the proof to Appendix F.3).
Lemma 5.
For every $L\in\mathsf{BQP}$-$\mathsf{complete}$ there exists an efficiently samplable distribution $\mathcal{D}_{L}$ such that $(L,\mathcal{D}_{L})\not\in\mathsf{HeurBPP}/\mathsf{samp}$, unless $\mathsf{DLP}\in\mathsf{BPP}$.
@@ 5 Connections to other works on (quantum) learning tasks
In this section we discuss other topics of relevance. First, in Section 5.1, we discuss the implications and limitations of the milestone work of Huang et al. [HKT+22b] on establishing learning separations from physical systems. Next, in Section 5.2, we discuss how having access to data radically enhances what can be efficiently evaluated by discussing the example of evaluating parameterized quantum circuits. Afterwards, in Section 5.3, we discuss how two physically-motivated problems (i.e., Hamiltonian learning, and identifying order parameters for phases of matter) fit in the PAC learning setting where the learner is constrained to output hypotheses from a fixed hypothesis class.
@@ 5.1 Provably efficient machine learning with classical shadows
In the milestone work of Huang et al. [HKT+22b], the authors design classical machine learning methods (in part built around the classical shadow paradigm) that can efficiently learn quantum many-body problems. One of the problems studied in [HKT+22b] is that of predicting ground states of Hamiltonian. More precisely, for a family of Hamiltonians $H(x)$ with ground states $\rho_{H}(x)$, one wants to predict the expectation value of some observable $O$ when measured on $\rho_{H}(x)$. That is, one wants to efficiently learn to evaluate the function
@@
One of the main things that [HKT+22b] show is that given a polynomial number of data points, one is able to efficiently evaluate the functions in Eq. (14) with a constant expected error under certain criteria. Recall that in Section 4.1 we argued that concepts based on physical systems can be used as a source of learning separations. Since these concepts are of a similar form as the functions described in Eq. (14), one might wonder how the results of Huang et al. relate.
Let us take a closer look at the requirements of the methods described in [HKT+22b]. Firstly, the Hamiltonians $H(x)$ must all be geometrically-local, and the observable $O$ must be a sum of polynomially many local observables $O=\sum_{i=1}^{L}O_{i}$ such that $\sum_{i=1}^{L}||O_{i}||$ is bounded by a constant. Additionally, the Hamiltonians $H(x)$ must all have a constant spectral gap (i.e., the difference between the smallest and the next smallest eigenvalue) and they must depend smoothly on $x$ (or more precisely, the average gradient of the function in Eq. (14) must be bounded by a constant). One might wonder what will happen if we relax the above requirements, while simultaneously maintaining the fact that a quantum computer would still be able to evaluate the function in Eq. (14) (and hence build a learning separation around it based on Theorem 9).
Two possible relaxations of the requirements are the absence of a constant spectral gap (while maintaining an inverse polynomial spectral gap) and a reduced smoothness dependency of the Hamiltonian family on $x$ (i.e., compared to what is required for the methods of [HKT+22b]). It turns out that if one relaxes these requirements, then under cryptographic assumptions the methods proposed by Huang et al. are no longer capable of evaluating the function in Eq. (14) with constant expected error. More precisely, any classical machine learning method that would still be able to evaluate the function in Eq. (14) up to constant expected error under the relaxed assumptions would be able to solve $\mathsf{DLP}$ in $\mathsf{P/poly}$, which contradicts certain cryptographic assumptions. We provide a formal statement of this in the following theorem, the proof of which is deferred to Appendix G.
@@
In conclusion, Theorem 10 shows that any method similar to that of [HKT+22b] cannot learn to predict ground state properties of certain physical systems discussed in Section 4.1. Moreover, there are a few subtle differences between the setup of [HKT+22b] and the one discussed in this paper. Firstly, the classical shadow paradigm uses data that is different from the PAC learning setting (i.e., the data does not correspond to evaluations of the function it aims to predict). This distinction in setup makes the approach of [HKT+22b] more versatile, as their data can be utilized to evaluate multiple different observables (moreover, their methods also work in the PAC setting). Secondly, the functions $f_{H,O}$ in Eq. (14) are real-valued, which differs from our paper where we investigate functions that map onto a discrete label space. It is possible to address this difference by applying a threshold function to $f_{H,O}$ after it is learned. However, this thresholding introduces a mismatch in the types of data, as it would involve using real-valued data to learn a function with discrete values (which is clearly different from the PAC setting).